OpenStudy (anonymous):

Extract the root: (x + 5)^2 = 10x + 74 explain step by step please

OpenStudy (anonymous):

(x + 5)^2 = x^2+10 x+25 now substituting: x^2+10 x+25 = 10x + 74 x^2+25 = + 74 x^2 = 49 x = +- 7

OpenStudy (anonymous):

I like Diogo answer. Summarizing the steps. 1. expand (multiply out) the squared term so (x+5) * (x+5) = X^2 + 10x + 25 2. so now we have x^2+10 x+25 = 10x + 74 3. gather like terms (move similar stuff all to one side): subtract 10x from both sides 4. now we have x^2 + 10x -10x + 25 = 10x -10x +74 or x^2 + 25 = 74 5. now subtract 25 from both sides: x^2 + 25 -25 = 74 - 25 6. leaves you with x^2 = 49 7. take square root so then x = \[\pm\]7